The Zeller Jewelry Company sold gold rings for $50 each and silver rings for $35 each. Total sales were $2,400. If customers bought twice as many silver as gold rings, how many of each did the store sell? Find the dollar sales of each.
Answer by richwmiller(17219)

The Zeller Jewelry Company sold gold rings for $50 each and silver rings for $35 each. Total sales were $2,400. If customers bought twice as many silver as gold rings, how many of each did the store sell? Find the dollar sales of each.
we need to set up two equations.
gold rings=g
silver rings=s
the first in order of information given
50g+35s=2400
2g=s
now wee substitute
50g+35*2g=2400
50g+70g=2400
120g=2400
g=20
s=40
But we are not finished yet. The questions doesn't ask for the number of silver and gold rings sold but the dollar amounts of the sales. We would have done this anyways to check our answers.

50*20=$1000 in gold rings
35*40=$1400 in silver rings
check
1000+1400=2400